# ParamANN: A Neural Network to Estimate Cosmological Parameters for $\Lambda$CDM Universe Using Hubble Measurements

### v2

@@ 3 Methodology
In section 3.1, we discuss the $H(z)$ measurements obtained by following the DA technique (available in the range $0.07\leq z\leq 1.965$). Then, in section 3.2 we describe the procedure to simulate the mock values of $H(z)$ and in section 3.3 we discuss the noise addition to these mock $H(z)$ data. Thereafter, in section 3.4 we provide the descriptions of the deep learning of the neural network used in our analysis. In section 3.5, we describe the MCMC approach which is performed to compare our AI technique.
@@ 3.2 Signal model
3.3 Noise model
We incorporate correlated Gaussian noises in the simulated $H(z)$ data for accounting the inherent noises of the Hubble measurements. We generate full covariance matrix corresponding to the Hubble measurements following the procedure of Moresco et al. (2020) [71]. In figure 1, we show this full covariance matrix which is used to generate the correlated Gaussian noises. We utilize the openly available code [^6] https://gitlab.com/mmoresco/CCcovariance [71] to estimate the covariances between Hubble parameters measured by Moresco et al. (2012, 2016) [66, 67] and Moresco (2015) [70]. Moreover, following the suggestion made by Moresco et al. (2020) [71], we consider the bias calculations for the systematic contributions due to initial mass function (IMF) and stellar population synthesis (SPS) model (odd one out) for the estimation of the covariances corresponding to these Hubble measurements [66, 67, 70]. We refer to the literature [66, 67, 70, 71] for the details about the systematic contributions in the Hubble parameters measured by the DA technique. For the rest [63, 64, 65, 68, 69] of observed Hubble data, we utilize only corresponding variances in the diagonal of the full covariance matrix. We consider the noise distributions around 00 mean, which are bounded by the full covariance matrix (shown in figure 1) corresponding to $31$ Hubble measurements. Using random.multivariate_normal [^7] https://numpy.org/doc/stable/reference/random/generated/numpy.random.multivariate_normal.html function of numpy library, we generate $1.2\times 10^{5}$ realizations of correlated Gaussian noises by using different seed values. Each of these noise realizations contains $31$ values corresponding to the observed redshift points (shown in table 1). We add these correlated noises to the simulated $H(z)$ data. These noise-included realizations of $H(z)$ are used as the input in ParamANN and the corresponding parameter values are provided as the targets in the output layer of ParamANN.
@@ 4.2.1 Hubble constant and density parameters
We note that, although the ParamANN is trained for generalized $\Lambda$CDM Universe by considering spatial curvature density, this trained ParamANN predicts spatially flat $\Lambda$CDM Universe by using DA Hubble measurements. Moreover, the parameters predicted by ParamANN are equivalent to the same estimated by MCMC analysis. In figure 6, we show $1\sigma$ and $2\sigma$ regions of 2D joint distribution of the parameters as well as the marginalized 1D marginalized probability densities of these parameters obtained by using the MCMC samples in GetDist. In figure 7, we demonstrate the estimated values of Hubble constant ($h_{0}$) and three density parameters ($\Omega_{0m}$, $\Omega_{0k}$ and $\Omega_{0\Lambda}$) with corresponding uncertainties provided by trained ParamANN and MCMC analysis. In the same figure, we also present the values of these parameters constrained by Planck collaboration [2]. We horizontally shift the parameter positions corresponding to the ParamANN and MCMC for the visual clarity of the overlapping points. Moreover, in this figure, we present the Hubble constant in 100 $\rm{kmMpc^{-1}s^{-1}}$ unit.
@@ 4.2.2 Hubble parameter curve
We obtain the Hubble parameter curves (equation (10)) using the parameter values (shown in the second and third columns of table 2) predicted by trained ParamANN as well as estimated by MCMC analysis. We also compute the same by using the parameter values (shown in the fourth column of table 2) constrained by Planck collaboration [2]. In figure 8, we show these three curves along with the data points (in the range $0.07\leq z\leq 1.965$) measured by the DA technique. Visually each of these three curves seems to fit the observations well. To test how these curves fit the observed points, we estimate the reduced-$\chi^{2}$ statistics as defined follow [^11] Although we have not done a $\chi^{2}$ fitting by ParamANN, we use the reduced-$\chi^{2}$ fit to estimate the goodness of the fits.,
@@ 5 Discussions and Conclusions
We create an ANN with one hidden layer containing 20 neurons. We call this ANN by a suitable name of ParamANN. We significantly train the ParamANN using $10^{5}$ samples of mock $H(z)$ values (which contain correlated noises convenient for observed Hubble data) as input and corresponding parameters as targets which are varied uniformly in their specified ranges, i.e. $\left[50,90\right]$ $\rm{kmMpc^{-1}s^{-1}}$, $\left[0.1,0.7\right]$ and $\left[0.3,0.9\right]$ for $H_{0}$, $\Omega_{0m}$ and $\Omega_{0\Lambda}$ respectively. We note that the mock values of $\Omega_{0k}$ is calculated by $1-\Omega_{0m}-\Omega_{0\Lambda}$. We use another $1.5\times 10^{4}$ samples of mock data to validate the training of ParamANN and $5\times 10^{3}$ samples to test the performance of trained ParamANN. We notice that the differences between targets and predictions for each parameter dominantly lie within three times uncertainties (predicted by ParamANN) corresponding to the predictions of test set. These results (shown in figure 5) of test set show the excellent agreement between targets and predictions of ParamANN (at least within the predicted uncertainties). Finally, we use the trained ParamANN for the predictions of the Hubble constant and the density parameters from the Hubble parameters measured by the DA technique in the redshift interval $0.07\leq z\leq 1.965$. We obtain $H_{0}=68.14\pm 3.96$ $\rm{kmMpc^{-1}s^{-1}}$, $\Omega_{0m}=0.3029\pm 0.1118$, $\Omega_{0k}=0.0708\pm 0.2527$ and $\Omega_{0\Lambda}=0.6258\pm 0.1689$. We note that the predicted density parameters show $0.07\sigma$, $0.28\sigma$ and $0.37\sigma$ deviations from Planck’s results for matter, curvature and vacuum respectively. Moreover, our predicted Hubble constant (showing $0.12\sigma$ deviation from Planck’s result) agrees well with Planck’s estimation of $H_{0}$.
@@
Acknowledgment
We acknowledge the use of open-source software library TensorFlow [^12] https://www.tensorflow.org/, python library numpy [^13] https://numpy.org/ and openly available code [^14] https://gitlab.com/mmoresco/CCcovariance given by Moresco. We thank Albin Joseph and Md Ishaque Khan for useful discussions associated with this work.

### v3

@@ 3 Methodology
In section 3.1, we discuss the $H(z)$ measurements obtained by following the DA technique (available in the range $0.07\leq z\leq 1.965$). Then, in section 3.2 we describe the procedure to simulate the mock values of $H(z)$ and
@@
in section 3.3 we discuss the noise addition to these mock $H(z)$ data. Thereafter, in section 3.4 we provide the descriptions of the deep learning of the neural network used in our analysis. In section 3.5, we describe the MCMC approach which is performed to compare our AI technique.
@@ 3.3 Noise model
We incorporate correlated Gaussian noises in the simulated $H(z)$ data to account the inherent noises of the Hubble measurements. We generate a full covariance matrix corresponding to the Hubble measurements following the procedure of Moresco et al(2020) [77]. In figure 1, we show this full covariance matrix which is used to generate the correlated Gaussian noises. We utilize the openly available code [^5] https://gitlab.com/mmoresco/CCcovariance [77] to estimate the covariances between Hubble parameters measured by Moresco et al(2012, 2016) [72, 73] and Moresco (2015) [76]. Moreover, following the suggestion made by Moresco et al(2020) [77], we consider the bias calculations for the systematic contributions due to the initial mass function (IMF) and stellar population synthesis (SPS) model (odd one out) for the estimation of the covariances corresponding to these Hubble measurements [72, 73, 76]. We refer to the literature [72, 73, 76, 77] for details about the systematic contributions in the Hubble parameters measured by the DA technique. For the rest [69, 70, 71, 74, 75] of observed Hubble data, we utilize only corresponding variances. In this figure 1, the diagonal values of the full covariance matrix represent the variances of the Hubble parameters measured in the redshift interval $0.07\leq z\leq 1.965$. Moreover, the off-diagonal region of this covariance matrix represents the covariances between these observed Hubble parameters. The values of variances and covariances are represented by the color bar in the figure 1. We note that the white pixels of this covariance matrix denote the redshift pairs for which no covariance information is available. We consider the noise distributions around 00 mean, which are bounded by the full covariance matrix (shown in figure 1) corresponding to $31$ Hubble measurements. Using random.multivariate_normal [^6] https://numpy.org/doc/stable/reference/random/generated/numpy.random.multivariate_normal.html function of numpy library, we generate $1.2\times 10^{5}$ realizations of correlated Gaussian noises by using different seed values. Each of these noise realizations contains $31$ values corresponding to the observed redshift points (shown in table 1). We add these correlated noises to the simulated $H(z)$ data. These noise-included realizations of $H(z)$ are used as the input in ParamANN and the corresponding parameter values are provided as the targets in the output layer of ParamANN.
@@ 4.2.1 Hubble constant and density parameters
We note that, although the ParamANN is trained for generalized $\Lambda$CDM Universe by considering spatial curvature density, this trained ParamANN predicts spatially flat $\Lambda$CDM Universe by using DA Hubble measurements. Moreover, the parameters predicted by ParamANN are equivalent to the same estimated by MCMC analysis. In figure 6, we show $1\sigma$ and $2\sigma$ regions of 2D joint distribution of the parameters as well as the marginalized 1D marginalized probability densities of these parameters obtained by using the MCMC samples in GetDist. In Appendix A, we include the Python code to obtain the figure 6 using GetDist. In figure 7, we demonstrate the estimated values of Hubble constant ($h_{0}$) and three density parameters ($\Omega_{0m}$, $\Omega_{0k}$ and $\Omega_{0\Lambda}$) with corresponding uncertainties provided by trained ParamANN and MCMC analysis. In the same figure, we also present the values of these parameters constrained by Planck collaboration [1]. We horizontally shift the parameter positions corresponding to the ParamANN and MCMC for the visual clarity of the overlapping points. Moreover, in this figure, we present the Hubble constant in 100 $\rm{kmMpc^{-1}s^{-1}}$ unit.
@@ 4.2.2 Hubble parameter curve
We obtain the Hubble parameter curves (equation (19)) using the parameter values (shown in the second and third columns of table 2) predicted by trained ParamANN as well as estimated by MCMC analysis. We also compute the same by using the parameter values (shown in the fourth column of table 2) constrained by Planck collaboration [1]. In figure 8, we show these three curves along with the data points (in the range $0.07\leq z\leq 1.965$) measured by the DA technique. We refer to Appendix B for the Python code of figure 8. Visually each of the three curves in figure 8 seems to fit the observations well. To test how these curves fit the observed points, we estimate the reduced-$\chi^{2}$ statistics as defined follow [^10] Although we have not done a $\chi^{2}$ fitting by ParamANN, we use the reduced-$\chi^{2}$ fit to estimate the goodness of the fits.,
@@ 5 Discussions and Conclusions
We create an ANN with one hidden layer containing 20 neurons. We call this ANN by a suitable name of ParamANN. We significantly train the ParamANN using $10^{5}$ samples of mock $H(z)$ values (which contain correlated noises convenient for observed Hubble data) as input and corresponding parameters as targets which are varied uniformly in their specified ranges, i.e. $\left[50,90\right]$ $\rm{kmMpc^{-1}s^{-1}}$, $\left[0.1,0.7\right]$ and $\left[0.3,0.9\right]$ for $H_{0}$, $\Omega_{0m}$ and $\Omega_{0\Lambda}$ respectively. We note that the mock values of $\Omega_{0k}$ is calculated by $1-\Omega_{0m}-\Omega_{0\Lambda}$. We use another $1.5\times 10^{4}$ samples of mock data to validate the training of ParamANN and $5\times 10^{3}$ samples to test the performance of trained ParamANN. We notice that the differences between targets and predictions for each parameter dominantly lie within three times uncertainties (predicted by ParamANN) corresponding to the predictions of test set. These results (shown in figure 5) of test set show the excellent agreement between targets and predictions of ParamANN (at least within the predicted uncertainties). Finally, we use the trained ParamANN for the predictions of the Hubble constant and the density parameters from the Hubble parameters measured by the DA technique in the redshift interval $0.07\leq z\leq 1.965$. We obtain $H_{0}=68.14\pm 3.96$ $\rm{kmMpc^{-1}s^{-1}}$, $\Omega_{0m}=0.3029\pm 0.1118$, $\Omega_{0k}=0.0708\pm 0.2527$ and $\Omega_{0\Lambda}=0.6258\pm 0.1689$. We note that the predicted density parameters show $0.07\sigma$, $0.28\sigma$ and $0.37\sigma$ deviations from Planck’s results for matter, curvature and vacuum respectively. Moreover, our predicted Hubble constant (showing $0.12\sigma$ deviation from Planck’s result) agrees well with Planck’s estimation of $H_{0}$. Furthermore, we obtain the Hubble parameter curves (figure 8) by using the estimated parameter values for each of ParamANN, MCMC and Planck collaboration [1]. We notice that these three curves excellently agree to each other and fit well to the observed Hubble data.
@@
We acknowledge the use of open-source software library TensorFlow [^11] https://www.tensorflow.org/, Python libraries numpy [^12] https://numpy.org/, GetDist [^13] https://getdist.readthedocs.io/en/latest/index.html and openly available code [^14] https://gitlab.com/mmoresco/CCcovariance given by Moresco. We thank Albin Joseph and Md Ishaque Khan for useful discussions associated with this work.
Data availability statement
The data can not be made publicly available upon publication because no suitable repository exists for hosting data in this field of study. The authors will share the data of this work after receiving a reasonable request.
Appendix A
We use the following Python code to obtain the figure 6.
  import getdist
from getdist import plots
import matplotlib.pyplot as plt
%matplotlib inline
plt.style.use('classic')
samples = getdist.MCSamples(samples=samples_array,
[…]
                  settings={'smooth_scale_1D':0.75,'smooth_scale_2D':0.75})
g = plots.get_subplot_plotter()
g.settings.axes_fontsize=20
g.settings.axes_labelsize=25
g.triangle_plot(samples,filled=True,line_args={'lw':1.5,'ls':'-',
                           'color':'teal'},contour_colors=['teal'])
ax1 = g.subplots[1, 0]
ax2 = g.subplots[2, 0]
ax3 = g.subplots[3, 0]
ax4 = g.subplots[3, 1]
ax5 = g.subplots[3, 2]
ax6 = g.subplots[3, 3]
g.rotate_yticklabels(ax=4,rotation=90,labelsize=None)
g.rotate_yticklabels(ax=8,rotation=90,labelsize=None)
g.rotate_yticklabels(ax=12,rotation=90,labelsize=None)
ax1.set_yticks([0.2,0.4,0.6],labels=[r'\$0.2\$',r'\$0.4\$',r'\$0.6\$'])
ax2.set_yticks([0.4,0.6,0.8],labels=[r'\$0.4\$',r'\$0.6\$',r'\$0.8\$'])
ax3.set_yticks([-0.5,0,0.5],labels=[r'-\$0.5\$',r'\$0\$',r'\$0.5\$'])
ax3.set_xticks([60,70,80],labels=[r'\$60\$',r'\$70\$',r'\$80\$'])
ax4.set_xticks([0.2,0.4,0.6],labels=[r'\$0.2\$',r'\$0.4\$',r'\$0.6\$'])
ax5.set_xticks([0.3,0.6,0.9],labels=[r'\$0.3\$',r'\$0.6\$',r'\$0.9\$'])
ax6.set_xticks([-0.5,0,0.5],labels=[r'-\$0.5\$',r'\$0\$',r'\$0.5\$'])
g.export('filename')
  where samples_array denotes the array of $6880$ samples of four parameters and filename defines the name of file to save the figure.
Appendix B
The following is the Python code to obtain the figure 8.
  import numpy as np
data = np.loadtxt('hubble.csv', delimiter=',',usecols=(0,1,2))
z = data[:,0]
Hz = data[:,1]
Hz_std = data[:,2]
import matplotlib
import matplotlib.pyplot as plt
import matplotlib.gridspec as gridspec
%matplotlib inline
plt.style.use('classic')
ann_Hz = 68.14*np.sqrt((0.3029*((1+z)**3))+(0.0708*((1+z)**2))+0.6258)
mcmc_Hz = 68.10*np.sqrt((0.2926*((1+z)**3))+(0.0626*((1+z)**2))+0.6416)
planck_Hz = 67.66*np.sqrt((0.3111*((1+z)**3))+(0.001*((1+z)**2))+0.6889)
fig = plt.figure(figsize=(10,7))
plt.errorbar(z,Hz,yerr=Hz_std,fmt='r*',ecolor='c',mec='r', markersize=8,
[…]
plt.plot(z,ann_Hz,'-',color='xkcd:green',linewidth=1,
[…]
plt.plot(z,mcmc_Hz,'-',color='xkcd:pumpkin',linewidth=1,
[…]
plt.plot(z,planck_Hz,'-',color='xkcd:blue',linewidth=1,
[…]
plt.grid(alpha=0.3)
plt.xticks([0.07,0.5,1,1.5,2],
                      labels=[r'\$0.07\$',r'\$0.5\$',r'\$1\$',r'\$1.5\$',r'\$2\$'])
plt.yticks([0,50,100,150,200,250],
                      labels=[r'\$0\$',r'\$50\$',r'\$100\$',r'\$150\$',r'\$200\$',r'\$250\$'])
plt.tick_params(axis='x',labelsize=20,pad=0.5)
plt.tick_params(axis='y',labelsize=20,pad=0,rotation=90)
plt.xlabel(r'\$z\$',fontsize=25,labelpad=-5)
plt.ylabel(r'\$H(z)\$',fontsize=25,labelpad=0)
plt.legend(loc='lower right',fontsize=20,numpoints=1,handlelength=1.0)
plt.savefig('filename',dpi=200,bbox_inches='tight',pad_inches=0.02)
  where three columns of hubble.csv data file contain observed redshifts (dimensionless), Hubble parameters and corresponding standard deviation in $\rm{kmMpc^{-1}s^{-1}}$ unit respectively. Moreover, filename represents the name of file to save the figure.